# Statistical analysis of geoinformation data for increasing railway safety

### v1

@@ Abstract
The impact of rail transport on the environment is one of the crucial factors for the sustainable development of this form of mass transport. We present a data-driven analysis of wild animal railway accidents in the region of southern Poland, a step to create the train driver warning system. We built our method by harnessing the Bayesian approach to the statistical analysis of information about the geolocation of the accidents. The implementation of the proposed a model does not require advanced knowledge of data mining and can be applied even in less developed railway systems with small IT support. Furthermore, we have discovered unusual patterns of accidents while considering the number of trains and their speed and time at particular geographical location of the railway network. We test the developed approach using data from southern Poland, compromising wildlife habitats and one of the most urbanised regions in Central Europe.
@@
  • easily implementable Bayesian model of accidents prediction was created,
@@
  • issues of wildlife ecology, and resilient and save rail transport were addressed,
@@ 1 Introduction
The demand for passenger transport is related to the quality of service and, in particular, to its resilience. It is very easy to lose passengers due to divergence from the planned schedule. Even a delay of several minutes causes some dissatisfaction of passengers, see Fig. 5 of [2]. Methodologically, disturbances in public transport may be divided into different categories according to dimension related to planning, probability, impact, time of occurrence, duration time, scope of location, frequency, and miscellaneous referring to the up-to-date review of disturbances in passenger traffic [3].
Collisions with wild animals are important examples of disturbances in passenger traffic that affect passenger satisfaction, particularly in the context of the competitiveness of public transport compared to private means of transport [4]. What is also important here is the fact that collisions with wild animals cause disturbances that propagate along the rail network, causing passengers dissatisfaction at various locations. This is not a desired situation, as the railways provide one of the most promising means of transportation from an environmental point of view. Moreover, reducing the number of collisions with wild animals has a positive effect on the environment itself, saving animals and the local heritage tied to them.
Aside from the passenger satisfaction issue, such collisions also have a negative impact on railway operation [5]. This includes the costs of the repair of the rolling stock, cleaning, etc. Besides, it is important to note that the construction of new types of rolling stock is more vulnerable to collision with animals than the old ones [6]. Furthermore, currently, trains move faster, but they are quieter [7] than in past years. This may reduce the effect of scaring away animals, especially in combination with particular landscape formations [8].
@@ 2.1 Review of predictive models
Starting from [10] predictive models have full meaning applications in the prediction of collisions with wild animals. The problem of trains colliding with wild animals appears all over the world [11]. The specifics of this problem are related to particular regions and spaces of animals leaving there; compare Asia [12] [13], North America [14, 15], South America [16], and Europe [17, 18, 19]. There can be drown the general conclusion from the analysis of such events – there appear hot-spots – places where the intensity of collisions is high. The challenge is to localise spatial and time coordinates in such places. This is the first step to undertake countermeasures to mitigate disturbances due to such accidents. In particular, these accidents are usually analysed with regard to the density of animals in the population, as well as the density of railway traffic. This leads to the spatial patterns approach, used in various analysis across the world. A prominent example of this approach is the analysis of accidents with elephants in India [20]. Another model demonstrates the link between kangaroo-train collisions, train speed, and the coincidence of periods of high train and kangaroo activity in Australia [21].
@@
Let us now concentrate on Europe, the region for which our analysis has been performed. The spatial patterns approach has been applied to analyse the collision of trains with moose (Alces alces) [22] in Norway. In a similar priori mentioned, hot-spots of animal collisions have been located in Romania and Sweden [23, 24]. Our analysis was performed for the south region of Poland, that is, close to Czechia, for which the analysis of the hot-spots has been identified using the KDE + (Kernel Density Estimation +) method [1]. Mentioned KDE+ extends the standard KDE (Kernel Density Estimation) by adding repeated random simulations using the Monte Carlo approach. In particular, the impact of various parameters on collisions has been analysed for the Czech railway network. The neighbourhood of the railway line has been parameterised, e.g. by the distance to the closest forest area, presence of water reservoirs and urban areas. Further parameters of the railway line have been considered, including maximum speed, presence of overhead electric wires, and track geometry [25].
@@
In Poland, analysis of trains collisions with wild animals for the time period 2012 - 2015 supplied by the national network operator (PKP PLK S.A.) was discussed in [26]. There, the spatial and time (day-time and year-time) distribution of collisions has been analysed and compared with the patterns of the wild life of various animals. For data analysis, generalised additive mixed models (GAMMs) were applied. Spatial analyses with open-source Geographic Information System (GIS) software were performed for chosen railway lines in the Greater Poland region, in particular on railway lines close to Poznań for 10 years (i.e., 2007 - 2017). In particular, spaces such as wild boar, red deer, roe deer, and lands such as forests, waters, developed areas, and agricultural land were analysed. The conclusion was that the analysis of collisions of trains with wild animals has to be performed separately for each space, as their behaviour is different for various lands [27].
@@
From the introduction and literature review we can conclude, that the analysis of time and spatial localisation of collisions with wild animals has been performed in Poland and other countries around the world, also using GIS tools. The results of the analysis demonstrate that there appears hot-spots, i.e. places where collisions are frequent. Furthermore, statistical analysis of basic features, such as time, particular spice ecological environment, speed of trains, etc. has been performed. This was tied to the possible implementation and analysis of acoustic animal warning devices.
@@ 2.2 Case-specific factors
One should note that in Poland the driver is obliged to stop the train immediately after the collision with the obstacle. Then the technical check of the rolling stock has to be performed. Frequently encountered faults are the violation of the vehicle sheathing, vehicle scrapers, damage to the sandblaster‘s wires, and in more serious cases, which may already have a significant impact on safety, these are failures of the braking systems, which may even result in the need to tow the vehicle. All of these problems add up to the delay and disturbance of the train. It is important to note that the driver has a range of tools that can be used to prevent a collision. These are: ringing the horn or applying emergency brakes [28], but there also exist some vehicle-based warning systems [29]. Hence, we believe that automatic warning before entering the zone where collisions with wild animals are possible would be the important tool in mitigating the problem of such collisions. To find such zones in time and space, we propose the systematic approach, based on the analysis of the locations and times of events.
@@ 3 Bayesian model of warning
In this section, we provide the details of the Bayesian model of marking line segments with warning. The model is based on past data on trains‘ collisions with animals. We start with an initial discussion of the underlying assumptions. Next, we provide a detailed description of the procedure, which can be implemented using any system that supports elementary statistical procedures.
@@ 3.2 Model details
The goal of our model is to take into consideration the spatial and temporal assessment of accidents. For this reason, we follow the recent approaches to train accidents with wild animals [20, 21]. In [20] trains collisions with elephants for years 1990-2018 in Indian railways were analysed. The spatial and temporal profiles of the elephant wildlife reservoirs were then estimated. In [21] the seven-year period 2009-2015 of data with collisions of trains with kangaroos was examined. There, the particular model of the temporal and spatial activity pattern of kangaroos is developed. Following these two papers, we also develop the risk model and predict the collision risk. However, in contrast to the models proposed in [20] and [21], our method is dedicated to the case where no particular data on activity patterns of animals (e.g. reservoirs, migrations) are available. Hence, we estimate all parameters of the model from the raw data of train animal collisions.
@@
  • $\mu(\tau)$ – the expected number of accidents in particular day, in month denoted by $\tau$; formally $\mu(\tau)=\sum_{k}kp(k,\tau)$ where $k$ numerates number of accidents in one day, and $p(k,\tau)$ is probability that $k$ accidents would occur in particular day in month $\tau$; however in our case even if $k>1$ various accidents occurs usually at different time intervals and multiple daily accidents are assumed to be independent;
@@
One should note that in our model, the independence of spatial and time profiles of accidents is assumed. This would hold in general as long as the migration paths of the animals did not have significant seasonal variations. We do not expect this in our regional approach. However, one should consider this point in the case of adapting the method to the particular case. Thanks to the assumption of independence of spatial and time profiles, we split our model into the spatial part
@@
Taking into account the above considerations, the probability that the particular train on the particular day takes part in the accident reads
@@ 4.1 Data description
The above-mentioned railway operator uses light electric multiple units (EMUs) that may be vulnerable to collisions with animals. Data on animal accidents are collected from drivers reports. Drivers are obliged, by the operator‘s regulations, to report any distortions and unusual events during the service of the train. Each such accident is annotated by the following set of parameters: date (day, month, year), time, number of line, kilometre post of the line, and the species. Data cover $3$ years in the period from 2020 to 2022. Additionally, we used a set of testing data from the first 4 months of 2023. From the 2020-2022 period, the profile of accidents has been demonstrated in Fig. 3 (time profiles) and Fig. 4 (spatial profiles).
@@ 4.2 Profile estimation
Here, $N_{\tau}$ is the number of accidents that appear in a month $\tau$, and $T/\tau_{\text{total}}$ is (approximated) number of days in particular month from data (for example if $T=365$ then $T/\tau_{\text{total}}\approx 30$).
@@
Let us now take advantage of the priori discussed assumption that time and spatial profiles are independent. Then, we can estimate $p(l)$ as
@@
where $\alpha(t,\Delta t)$ is the temporal profile taking into account pick hours and off pick hours and night break
@@
It is derived from the observation that in pick hours, i.e. 6 a.m. - 9 a.m. and 3 p.m. - 5 p.m., $40\%$ of trains are served. Further, it is assumed that in night hours, 12 a.m. - 3 p.m., only $5\%$ of trains are served. The rest of the trains are served during pick off hours. Finally, one can use the formula provided in Eq. (4) to compute the probability of the accident of the particular train in the particular day given $\tau$, $t$, $\Delta t$, $l$, $x$, $\Delta x$.
@@ 4.3 Illustrative example
    Using Eqs. (10), (11) at the off pick-up hour, we have $m(\tau,t,\Delta t,l,x,\Delta x)=131\cdot 0.55/14\approx 5.15$.
@@ 4.4 Evaluation results
The large number of accidents above $80$ km post-line 139 appeared regardless of low speed in this fragment of line. This observation is contrary to the observation in [12] for elephants, where train speed was recognised as one of the main factors in animal accidents. At this part of line $139$ train speed is limited (approximately $50$ km/h; see lower panel of Fig. 6). To evaluate this important observation, in Fig. 7 we demonstrate the relation between train speed and the number of accidents per train for chosen lines. What is important is that we have not observed the positive correlation between these two quantities. This is most likely because many accidents occur in low-speed forest areas, where railway traffic is not particularly dense. Therefore, in this area, animals are not scared away from the track by fast moving trains. This important observation should be tested on more of railway systems, where accidents with animals like roe deer or wild boar are recorded.

### v2

@@ Abstract
The impact of rail transport on the environment is one of the crucial factors for the sustainable development of this form of mass transport. We present a data-driven analysis of wild animal railway accidents in the region of southern Poland, a step to create the train driver warning system. We built our method by harnessing the Bayesian approach to the statistical analysis of information about the geolocation of the accidents. The implementation of the proposed model does not require advanced knowledge of data mining and can be applied even in less developed railway systems with small IT support. Furthermore, we have discovered unusual patterns of accidents while considering the number of trains and their speed and time at particular geographical locations of the railway network. We test the developed approach using data from southern Poland, compromising wildlife habitats and one of the most urbanised regions in Central Europe, based on this we conclude that our model is best suited to railway lines that pass through varying types of landscape.
@@
  • easily implementable Bayesian model of accident prediction was created,
@@
  • issues of wildlife ecology, and resilient and safe rail transport were addressed,
@@ 1 Introduction
The demand for passenger transport is related to the quality of service and, in particular, to its resilience. It is very easy to lose passengers due to divergence from the planned schedule. Even a delay of several minutes causes some dissatisfaction among passengers, see Fig. 5 of [2]. Methodologically, disturbances in public transport may be divided into different categories according to dimensions related to planning, probability, impact, time of occurrence, duration time, scope of location, frequency, and miscellaneous referring to the up-to-date review of disturbances in passenger traffic [3].
Collisions with wild animals are important examples of disturbances in passenger traffic that affect passenger satisfaction, particularly in the context of the competitiveness of public transport compared to private means of transport [4]. What is also important here is the fact that collisions with wild animals cause disturbances that propagate along the rail network, causing passenger dissatisfaction at various locations. This is not a desired situation, as the railways provide one of the most promising means of transportation from an environmental point of view. Moreover, reducing the number of collisions with wild animals has a positive effect on the environment itself, saving animals and the local heritage tied to them.
Aside from the passenger satisfaction issue, such collisions also have a negative impact on railway operation [5]. This includes the costs of the repair of the rolling stock, cleaning, etc. Besides, it is important to note that the construction of new types of rolling stock is more vulnerable to collision with animals than the old ones [6]. Furthermore, currently, trains move faster, but they are quieter [7] than in past years. This may reduce the effect of scaring away animals, especially in combination with particular landscape formations [8]. The overall socio-economic costs of train collisions with wildlife in Sweden have recently been estimated at 100,000,000–150,000,000 Euros per year [9]. This is similar to the costs estimated for wildlife-vehicle collisions on roads (250,000,000 Euros per year) [10], even though railways comprise less than 2% of the national road network in Sweden. Such an observation provides the sound motivation for this research. Concerning Poland and referring to the Intercity operator (PKP Intercity, eng. Polish State Railways Intercity), the cost of repair of rolling stock after collisions with wild animals in 2014 was estimated at the value above 100,000,000 Euros [11].
@@ 2.1 Review of predictive models
Starting from [13] predictive models have full meaning applications in the prediction of collisions with wild animals. The problem of trains colliding with wild animals appears all over the world [14]. The specifics of this problem are related to particular regions and spaces of animals leaving there; compare Asia [15] [16], North America [17, 18], South America [19], and Europe [20, 21, 22]. There can be drown the general conclusion from the analysis of such events – there appear hot-spots – places where the intensity of collisions is high. The challenge is to localise spatial and temporal coordinates in such places. This is the first step to undertake countermeasures to mitigate disturbances due to such accidents. In particular, these accidents are usually analysed with regard to the density of animals in the population, as well as the density of railway traffic. This leads to the spatial patterns approach, used in various analysis across the world. A prominent example of this approach is the analysis of accidents with elephants in India [23]. Another model demonstrates the link between kangaroo-train collisions, train speed, and the coincidence of periods of high train and kangaroo activity in Australia [24].
@@
Let us now concentrate on Europe, the region for which our analysis has been performed. The spatial patterns approach has been applied to analyse the collisions of train with moose (Alces alces) [25] in Norway. In a similar priori mentioned, hot-spots of animal collisions have been located in Romania and Sweden [10, 26]. Our analysis was performed for the south region of Poland, that is, close to Czechia, for which the analysis of the hot-spots has been identified using the KDE + (Kernel Density Estimation +) method [1]. Mentioned KDE+ extends the standard KDE (Kernel Density Estimation) by adding repeated random simulations using the Monte Carlo approach [27]. In particular, the impact of various parameters on collisions has been analysed for the Czech railway network. The neighbourhood of the railway line has been parameterised, e.g. by the distance to the closest forest area, the presence of water reservoirs, and urban areas. Further parameters of the railway line have been considered, including maximum speed, presence of overhead electric wires, and track geometry [28].
article animal region
use external data
predictive method
[29]
Tree frog
France
Yes (habitats)
landscape
graph construction
[15] Elephant India Yes (eleph. signs)
Kernel/ point density
[18]
Bear, carnivore
ungulate
Canada
Yes (proximity to water,
amount of water)
Regression
multi parameters
[21] Tortoise Romania Yes (habitats)
Factor model
tied to habitants
[22] Bat Poland
Yes (bat routine
and habitats observation)
General linear mixed
model (GLMM)
[23] Elephant India
Yes (spatial relationship
elephant density)
Geographically
weighted regression
[24] Kangaroo Australia
Yes (kangaroo occurrence)
Regression model
[25] Moose Norway Yes (vegetation removal) Cost-benefit analysis
[10]
Moose, roe deer,
reindeer
Sweden
Yes (composition of the
surrounding landscape,
availability of forage,
preferred habitat)
Logistic regression
[26]
Brown bear, lynx,
wolf, red deer,
roe deer, wild boar
Romania
Yes (number of
individual at location)
Weighted factors
analysis
[1]
Roe deer, wild boar,
red deer, red fox,
fallow deer, European
badger, European hare,
muflon
Czechia
No
Kernel density
+ Monte Carlo
method derived for
animals cars collisions
[30]
Roe deer, red deer,
moose, wild boar
Poland No
Generalised additive
mixed models
(GAMMs) uses only
temporal data
[31]
Roe deer, wild boar,
red deer
Poland
Yes (presence of specific
land use function)
Spatial analyses
(QGIS, QChainge)
Table 1: Summarising and comparison of relevant studies on predictive models of rail animals collisions. Observe, that there are only two works that do not consider external data from the landscape.
@@
In Poland, analysis of train collisions with wild animals for the time period 2012 - 2015 supplied by the national network operator (PKP PLK S.A.) was discussed in [30]. There, the temporal (day-time and year-time) distribution of collisions has been analysed and compared with the patterns of the wildlife of various animals. For data analysis, generalised additive mixed models (GAMMs) were applied. Spatial analyses with open-source Geographic Information System (GIS) software were performed for chosen railway lines in the Greater Poland region, in particular on railway lines close to Poznań for 10 years (i.e., 2007 - 2017). In particular, spaces such as wild boar, red deer, roe deer, and lands such as forests, waters, developed areas, and agricultural land were analysed. The conclusion was that the analysis of collisions of trains with wild animals has to be performed separately for each space, as their behaviour is different for various lands [31].
@@
From the introduction and literature review we can conclude, that the analysis of temporal and spatial localisation of collisions with wild animals has been performed in Poland and other countries around the world, also using GIS tools. The results of the analysis demonstrate that there appears hot-spots, i.e. places where collisions are frequent. Furthermore, statistical analysis of basic features, such as time, particular spice ecological environment, speed of trains, etc. has been performed. This was tied to the possible implementation and analysis of acoustic animal warning devices.
he relevant studies on predictive models of train-animals collisions are summarised in Table 1. From this summary, one can conclude that most studies input external data (called also landscape data) such as landscape form, animal paths, animal habitat, etc. These data are problematic to acquire by the railway operator, making such a model less practical. We have found two works that do not consider such external data, the first [1] uses a Kernel density + Monte Carlo method that is more dedicated to road-animals collisions and uses only limited temporal relation of data, the second method [30] does not consider spatial relation of data, henceforth it cannot be applied in hot-spot detection. We recognise the scientific gap of the lack of systematic research on railway-dedicated collision hot-spot detection, given statistical data that the railway operators can easily acquire.
@@ 2.2 Case-specific factors
One should note that in Poland, the driver is obliged to stop the train immediately after the collision with the obstacle. Then the technical check of the rolling stock has to be performed. Frequently encountered faults are the violation of the vehicle sheathing, vehicle scrapers, damage to the sandblaster‘s wires, and in more serious cases, which may already have a significant impact on safety, these are failures of the braking systems, which may even result in the need to tow the vehicle. All of these problems add up to the delay and disturbance of the train. It is important to note that the driver has a range of tools that can be used to prevent a collision. These are: ringing the horn or applying emergency brakes [32], but there also exist some vehicle-based warning systems [33]. Hence, we believe that automatic warning before entering the zone where collisions with wild animals are possible would be the important tool in mitigating the problem of such collisions. To find such zones in time and space, we propose the systematic approach, based on the analysis of the locations and times of events.
@@ 3 Bayesian model of warning
In this section, we provide the details of the Bayesian model of marking line segments with warning. The model is based on past data on train collisions with animals. We start with an initial discussion of the underlying assumptions. Next, we provide a detailed description of the procedure, which can be implemented using any system that supports elementary statistical procedures.
@@ 3.2 Model details
The goal of our model is to take into consideration the spatial and temporal assessment of accidents. For this reason, we follow the recent approaches to train accidents with wild animals [23, 24]. In [23] train collisions with elephants for years 1990-2018 in Indian railways were analysed. The spatial and temporal profiles of the elephant wildlife reservoirs were then estimated. In [24] the seven-year period 2009-2015 of data with collisions of trains with kangaroos was examined. There, the particular model of the temporal and spatial activity pattern of kangaroos is developed. Following these two papers, we also develop the risk model and predict the collision risk. However, in contrast to the models proposed in [23] and [24], our method is dedicated to the case where no particular data on activity patterns of animals (e.g. reservoirs, migrations) are available. Hence, we estimate all parameters of the model from the raw data of train animal collisions.
@@
  • $\mu(\tau)$ – the expected number of accidents in particular day, in month denoted by $\tau$; formally $\mu(\tau)=\sum_{k}kp(k,\tau)$ where $k$ numerates number of accidents in one day, and $p(k,\tau)$ is probability that $k$ accidents would occur in particular day in month $\tau$; however in our case even if $k>1$ various accidents usually occurs at different time intervals and multiple daily accidents are assumed to be independent;
@@
One should note that in our model, the independence of spatial and temporal profiles of accidents is assumed. This would hold in general as long as the migration paths of the animals did not have significant seasonal variations. We do not expect this in our regional approach. However, one should consider this point in the case of adapting the method to the particular case. Thanks to the assumption of independence of spatial and temporal profiles, we split our model into the spatial part
@@
Considering the above considerations, the probability that the particular train on the particular day takes part in the accident reads
@@ 4.1 Data description
The above-mentioned railway operator uses light electric multiple units (EMUs) that may be vulnerable to collisions with animals. Data on animal accidents are collected from drivers reports. Drivers are obliged, by the operator‘s regulations, to report any distortions and unusual events during the service of the train. Each such accident is annotated by the following set of parameters: date (day, month, year), time, number of line, kilometre post of the line, and the species. Data cover $3$ years in the period from 2020 to 2022. Additionally, we used a set of testing data from the first 4 months of 2023. From the 2020-2022 period, the profile of accidents has been demonstrated in Fig. 3 (time profiles) Fig. 4 (detailed monthly profiles) and Fig. 5 (spatial profiles).
Figure 4: Monthly profiles of accidents involving wild animals over particular years. We observe similar profile for all years. To demonstrate the stability of model, one can compare 2020 4 and 2021 4 as in the first half of 2020 train traffic was reduced due to COVID-19 lockdown.
@@
In Table 2 we present monthly statistics of train collision data to account for the size of the data set and its variability on a yearly basis. Referring to Table 2 and to profiles in Fig. 4 we can conclude no meaningful differences between years. This demonstrates the robustness of the method, regardless of the COVID-10 lockdown in the first half of 2020.
yearly monthly
year total min. max. mean std
2020 264 14 33 22.0 5.67
2021 348 19 46 29.0 8.37
2022 265 12 39 22.1 8.39
Table 2: Monthly statistics of number of accidents. Observe, that yearly means are similar, which reflects the statistical stability of data.
@@ 4.2 Profile estimation
Here, $N_{\tau}$ is the number of accidents that appear in a month $\tau$, and $T/\tau_{\text{total}}$ is (approximated) the number of days in a particular month from data (for example if $T=365$ then $T/\tau_{\text{total}}\approx 30$).
@@
Let us now take advantage of the priori-discussed assumption that temporal and spatial profiles are independent. Then, we can estimate $p(l)$ as
@@
where $\alpha(t,\Delta t)$ is the temporal profile considering peak hours and off-peak hours and night break
@@
It is derived from the observation that in peak hours, i.e. 6 a.m. - 9 a.m. and 3 p.m. - 5 p.m., $40\%$ of trains are served. Further, it is assumed that in night hours, 12 a.m. - 3 p.m., only $5\%$ of trains are served. The rest of the trains are served during off-peak hours. Finally, one can use the formula provided in Eq. (4) to compute the probability of the accident of the particular train on a particular day given $\tau$, $t$, $\Delta t$, $l$, $x$, $\Delta x$.
@@ 4.3 Illustrative example
    Using Eqs. (10), (11) at the off-peak, we have $m(\tau,t,\Delta t,l,x,\Delta x)=131\cdot 0.55/14\approx 5.15$.
@@ 4.4 Evaluation results
The large number of accidents above $80$ km post-line 139 appeared regardless of low speed in this fragment of line. This observation is contrary to the observation in [15] for elephants, where train speed was recognised as one of the main factors in animal accidents. At this part of line $139$ train speed is limited (approximately $50$ km/h; see lower panel of Fig. 7). To evaluate this important observation, in Fig. 8 we demonstrate the relation between train speed and the number of accidents per train for chosen lines. What is important is that we have not observed the positive correlation between these two quantities. This is most likely because many accidents occur in low-speed forest areas, where railway traffic is not particularly dense. Therefore, in this area, animals are not scared away from the track by fast-moving trains. This important observation should be tested on more of railway systems, where accidents with animals like roe deer or wild boar are recorded.
@@
To present metrics of the model‘s performance, we opt for the accuracy measure computed as the $p_{\text{threshold}}$ dependent fraction of accidents from the testing set that fall into warnings computed via Procedure 1 from the training set. This accuracy is compared with the $p_{\text{threshold}}$ dependent percentage of warnings, namely: the actual number of warnings over a maximal possible number of warnings. The maximal possible number of warnings is the number of warnings at $p_{\text{threshold}}=0$ (indeed a little below 00 as we have a sharp inequality in Eq. (5)).
For the comprehensive analysis we compare our method with the one reviewed form [30] (regarded earlier as closest to our approach). In [30] only temporal (day-time and year-time) dependency of accidents was considered, disregarding the spatial dependency. Henceforth, our goal is to assess the impact of spatial dependency on the model‘s effectiveness. Henceforth, we use two different testing scenarios.
  1. 1.
    Bayesian, the spatial and temporal - ST, where we check whether each accident (recorded in the testing set) fits into the spatial and temporal warning window computed by Procedure 1 from the training set.
  2. 2.
    Derived from [30], only temporal- T, the whole line at a given time is considered to be under the warning if there appears the warning computed by Procedure 1, at any location on the line at this time.
Results of testing these scenarios, compared with random choice, are presented in Figure. 11 - the plot was performed for various $p_{\text{threshold}}$ values ranging from 00 to $0.01$.
Figure 11: Metrics of model effectiveness (measured as the percentage of accidents from testing data set that fits into warnings computed from the training set), for various lines and various $\Delta x$ parameter in Procedure 1. ST - original spatial and temporal testing scenario, T - temporal only testing scenario, where spatial dependency was not considered during testing. Observe, that for line number $1$ the ST testing scenario returns better results.
The results in fig. 11 (top panel) indicate that in the case of line number $1$ the spatial and temporal (ST) testing scenario returns better results, than the temporal (T) approach. This is the sound result as line number $1$ passes through variable landscapes (industrial, agricultural, and forests). This observation confirms the intuition that our approach is most effective for lines that pass varying types of landscape.
The results in Fig. 11 (bottom panel) indicate similar results of ST and T testing scenario for aggregated data from lines number $1$, $139$, and $140$. Furthermore, the TS scenario slightly outperforms the T one for particular $\Delta x$ settings (Fig. 11 (d)). From the train driver‘s point of view, the ST approach is more sound as it has the potential for hot-spot detection. This observation paves the path for the elaboration of new models on more precise hot-spot detection algorithms from the statistical data only. The important issue here is to tune model parameters such as spatial resolution $\Delta x$ in Procedure 1.